OpenStudy (anonymous):

Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal. My function is A(t)=39146(1+0.03/12)^12t, It would be very nice if someone could through this with me step by step @joquez14

OpenStudy (anonymous):

is this all the info

OpenStudy (anonymous):

@DevilInADress are you there

OpenStudy (anonymous):

No It's not. I will send it to you. Please give e a sec

OpenStudy (anonymous):

You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3%

OpenStudy (anonymous):

and do you know what the variables A(t) mean

OpenStudy (anonymous):

Is it the amount the car cost?

OpenStudy (anonymous):

yes and t is the months

OpenStudy (anonymous):

I'm not sure if it will help but it was said that the answer for this question will be negative. I am guessing because the decrease in months ?

OpenStudy (anonymous):

\[10,000=39,145(1+\frac{ .03 }{ 12 }^{12t}\] is your equation now can you solve for t

OpenStudy (anonymous):

Can you tell me what to do and I will go from there then check back to see if it is correct ?

OpenStudy (anonymous):

first you would divide 39,145 by both sides \[\frac{10,000}{39,145} = (1+\frac{.03}{12})^{12t}\]

OpenStudy (anonymous):

n\[.25546= (1+\frac{.03}{12})^{12t}\]

OpenStudy (anonymous):

do you know how to log

OpenStudy (anonymous):

No idea

OpenStudy (anonymous):

so you get an idea read this http://www.mathsisfun.com/algebra/logarithms.html

OpenStudy (anonymous):

so after you logged it your answer should be t= -4.605

OpenStudy (anonymous):

Would i just type this as my explanation ?

OpenStudy (anonymous):

yes just put t = -4.605 and show the steps i did and you should get it right

OpenStudy (anonymous):

Okay Thank You. Can you help me with one more ? pretty please?

OpenStudy (anonymous):

ill try